# subequation

### v1

@@ 00 $s$ $\Sigma_{u}^{-}$$\{0^{++},1^{+-}\}$ $H_{3}$
$\ell$ $n$ $M^{(0)}$ $\Gamma_{(i)}$ $\Gamma_{(ii)}$ $\Gamma_{(iii)}$ $\delta E_{(i)}$ $\delta E_{(ii)}$ $\delta E_{(iii)}$ $\Gamma^{s}_{(i)}$ $\Gamma^{s}_{(ii)}$ $\Gamma^{s}_{(iii)}$ $\delta E^{s}_{(i)}$ $\delta E^{s}_{(ii)}$ $\delta E^{s}_{(iii)}$
0 1 $4554$ $7$ $<1$ $5$ $58$ $31$ $37$ $19$ $4$ $14$ $45$ $27$ $30$
0 2 $4996$ $2$ $<1$ $1$ $33$ $15$ $20$ $5$ $<1$ $3$ $23$ $11$ $14$
0 3 $5388$ $<1$ $1$ $<1$ $21$ $8$ $12$ $1$ $<1$ $<1$ $14$ $5$ $8$
0 4 $5747$ $<1$ $1$ $<1$ $15$ $6$ $8$ $<1$ $<1$ $<1$ $9$ $3$ $5$
0 5 $6081$ $<1$ $2$ $<1$ $11$ $4$ $6$ $<1$ $<1$ $<1$ $6$ $2$ $4$
1 1 $4110$ $36$ $28$ $24$ $-29$ $-13$ $-17$ n/a n/a n/a $-9$ $-5$ $-5$
1 2 $4462$ $49$ $23$ $29$ $-9$ $-1$ $-5$ $20$ $11$ $13$ $-10$ $-4$ $-6$
1 3 $4765$ $6$ $<1$ $3$ $20$ $14$ $13$ $8$ $2$ $5$ $14$ $11$ $9$
1 4 $4814$ $28$ $14$ $17$ $-8$ $-1$ $-4$ $11$ $6$ $7$ $-6$ $-2$ $-4$
1 5 $5130$ $23$ $7$ $13$ $5$ $3$ $3$ $14$ $5$ $8$ $2$ $1$ $1$
2 1 $4321$ $24$ $22$ $16$ $-8$ $-3$ $-5$ $5$ $7$ $4$ $-6$ $-5$ $-4$
2 2 $4658$ $30$ $19$ $19$ $-4$ $<1$ $-2$ $14$ $10$ $9$ $-5$ $-2$ $-3$
2 3 $4965$ $10$ $4$ $6$ $9$ $8$ $6$ $8$ $3$ $5$ $5$ $5$ $3$
2 4 $5001$ $16$ $11$ $10$ $-4$ $<1$ $-2$ $6$ $5$ $4$ $-3$ $<1$ $-2$
2 5 $5297$ $20$ $9$ $12$ $2$ $2$ $1$ $12$ $6$ $7$ $<1$ $<1$ $<1$
@@
Table 4: Contributions of $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}$ to the hybrid charmonium widths and masses. All dimension-full entries are in MeV units. The angular momentum of the heavy meson-antimeson pair is $l^{\prime}=\ell$. The subscript with Roman numbers refers to the model potential used (see Fig (2)). The superscript $s$ indicates the that the quantities refer to the heavy-meson pairs with strangeness. Energy contributions labeled as $<1$ should be understood as the absolute value of the contribution being smaller than $1$ MeV. The third column corresponds to $M^{(0)}=E^{(0)}+2m_{Q}$.
$\ell$ $n$ $M^{(0)}$ $\Gamma_{(i)}$ $\Gamma_{(ii)}$ $\Gamma_{(iii)}$ $\delta E_{(i)}$ $\delta E_{(ii)}$ $\delta E_{(iii)}$ $\Gamma^{s}_{(i)}$ $\Gamma^{s}_{(ii)}$ $\Gamma^{s}_{(iii)}$ $\delta E^{s}_{(i)}$ $\delta E^{s}_{(ii)}$ $\delta E^{s}_{(iii)}$
0 1 $11077$ $2$ $5$ $<1$ $107$ $51$ $71$ $3$ $<1$ $6$ $87$ $46$ $62$
0 2 $11361$ $1$ $5$ $<1$ $67$ $28$ $42$ $4$ $<1$ $4$ $54$ $23$ $35$
0 3 $11616$ $<1$ $5$ $<1$ $46$ $17$ $28$ $2$ $<1$ $2$ $36$ $14$ $22$
0 4 $11851$ $1$ $4$ $<1$ $33$ $12$ $19$ $1$ $<1$ $<1$ $25$ $9$ $15$
0 5 $12071$ $1$ $4$ $<1$ $25$ $9$ $15$ $<1$ $<1$ $<1$ $19$ $6$ $11$
1 1 $10783$ $81$ $55$ $64$ $-59$ $-14$ $-28$ n/a n/a n/a $-15$ $-7$ $-10$
1 2 $10980$ $129$ $45$ $74$ $-21$ $<1$ $-8$ $36$ $18$ $25$ $-29$ $-9$ $-15$
1 3 $11176$ $71$ $18$ $41$ $34$ $15$ $20$ $58$ $18$ $35$ $15$ $9$ $9$
1 4 $11219$ $14$ $11$ $8$ $25$ $21$ $17$ $<1$ $<1$ $<1$ $26$ $18$ $18$
1 5 $11380$ $63$ $17$ $35$ $6$ $3$ $4$ $36$ $11$ $20$ $-3$ $<1$ $-2$
2 1 $10895$ $63$ $39$ $40$ $-19$ $-4$ $-10$ $4$ $6$ $4$ $-14$ $-8$ $-9$
2 2 $11092$ $87$ $40$ $52$ $-13$ $<1$ $-6$ $29$ $17$ $19$ $-17$ $-7$ $-10$
2 3 $11284$ $75$ $27$ $43$ $9$ $7$ $6$ $44$ $18$ $26$ $-2$ $1$ $<1$
2 4 $11346$ $3$ $4$ $2$ $20$ $18$ $13$ $<1$ $<1$ $<1$ $19$ $14$ $12$
2 5 $11480$ $64$ $23$ $37$ $<1$ $2$ $<1$ $32$ $13$ $19$ $-6$ $-1$ $-3$
@@
Table 5: Contributions of $B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)}$ to the hybrid bottomonium widths and masses. All dimension-full entries are in MeV units. The angular momentum of the heavy meson-antimeson pair is $l^{\prime}=\ell$. The subscript with Roman numbers refers to the model potential used (see Fig (2)). The superscript $s$ indicates the that the quantities refer to the heavy-meson pairs with strangeness. Energy contributions labeled as $<1$ should be understood as the absolute value of the contribution being smaller than $1$ MeV. The third column corresponds to $M^{(0)}=E^{(0)}+2m_{Q}$.
@@
We display our results for energy shifts caused by the heavy meson pairs to the hybrid masses and the decay widths of the hybrid states into the heavy meson pairs in tables 4 and 5 for charmonium and bottomonium hybrids, respectively. In the charmonium sector we find that the energy contributions are, with a few exceptions, below the $30$ MeV, thus being smaller than the $\mathcal{O}(1/m_{Q})$ spin-dependent operator contributions Brambilla:2018pyn ; Brambilla:2019jfi ; Soto:2023lbh and also smaller than the expected size of the spin-independent ones, which so far have not been computed. The energy contributions tend to decrease with increasing $n$ and $\ell$. The widths range between $1$ to $20$ MeV per channel [^5] Notice, that neutral and charged $D$ meson contributions are added up in columns 4 to 9 of table 4.. The larger widths are found in states closer to the heavy meson pair thresholds, with the values being very sensitive to the energy gap between the state and threshold. Adding up the widths for the two channels studied, we find values that range from slightly larger than the semi-inclusive widths for transitions to quarkonium Oncala:2017hop ; Brambilla:2022hhi to much smaller ones. The latter, however, correspond to higher lying hybrid charmonium states that can also decay to higher mass heavy-meson pairs not considered here.
@@
In the bottomonium sector, the energy contributions are larger than in the charmonium sector, with up to ten states having contributions larger than $30$ MeV. As a consequence for bottomonium hybrids the $B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)}$ contribution to the mass is significantly larger than the spin-dependent one. The widths are a factor of $2$ to $3$ larger than in the charmonium sector. The larger contributions in the bottomonium sector can be traced back to the fact that the coupling operator is not heavy-quark mass suppressed and the less extended hybrid bottomonium wave functions having a larger overlap with the potential than the more extended wave functions of hybrid charmonia. In contrast, semi-inclusive transition widths to quarkonium are smaller for bottomonium hybrids than for charmonium hybrids Oncala:2017hop ; TarrusCastella:2021pld ; Brambilla:2022hhi and consequently also smaller than decay widths to $B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)}$. As in the charmonium sector, the lowest mass $\ell=0$ states have the largest contributions to the mass, while the larger widths are found in the lowest lying $\ell=1$ states.
@@
It is perhaps unexpected that the widths for $\ell=0$ are smaller than those of other states of similar mass, while the energy shifts are larger. This can be explained as follows. The transition form factors into $S$-wave heavy meson pairs, $a_{n0;S}(k)$, have the largest peak, resulting in large values for the energy shifts produced by the heavy meson pairs, however, they decrease very rapidly to zero as $k$ increases. As the $\ell=0$ states are far away from the thresholds the values of $p_{r}$ are large enough for the widths to correspond to the region where the form factors are small. Moreover, as the energy contribution is overall positive it tends to increase the value of hybrid mass and, correspondingly, of the value of $p_{r}$, further decreasing the value of the width.
@@
In table 6 we show the spectra of the pure $\Pi_{u}$ states that do not couple with the two lowest lying thresholds at leading order. The lowest mass states, $n=1$, $\ell=1$, in both sectors are interesting, since they should have very suppressed decays into $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ and are below any other heavy meson pair threshold. Furthermore, two of the spin triplet states of corresponding multiplet have explicitly exotic $j^{pc}$. If they could be measured, identifying these states as hybrids should be straightforward due to these two very distinct characteristics.
$\ell$ $n$ $M^{(0)}_{c}$ $M^{(0)}_{b}$
1 1 $4238$ $10840$
1 2 $4610$ $11059$
1 3 $4954$ $11267$
1 4 $5277$ $11466$
1 5 $5583$ $11657$
2 1 $4429$ $10945$
2 2 $4786$ $11160$
2 3 $5118$ $11364$
2 4 $5433$ $11560$
2 5 $5732$ $11748$
@@
Table 6: Masses of the pure $\Pi_{u}$ hybrid quarkonium states that do not couple to $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ corresponding to the $H_{2}(\ell=1)$ and $H_{5}(\ell=2)$ multiplets of table 3. The masses are obtained as $M^{(0)}=E^{(0)}+2m_{Q}$ and are given in MeV units.
@@
Figure 3: Comparison of the mass and widths of hybrid charmonium states in table 4 with experimental states above threshold listed in the PDG Workman:2022ynf . The masses and widths of the hybrid states are obtained as the average of the results corresponding to the three model potentials in Fig. 2. The bars correspond to the sum in quadrature of the standard deviation of the results for the three models and the size of higher order corrections $\mathcal{O}(\Lambda_{\rm QCD}/m_{c})$.

### v2

@@ 00 $s$ $\Sigma_{u}^{-}$$\{0^{++},1^{+-}\}$ $H_{3}$
$n\,l_{\ell}$ $M_{c}^{(0)}$ $\Gamma_{(i)}$ $\Gamma_{(ii)}$ $\Gamma_{(iii)}$ $\delta E_{(i)}$ $\delta E_{(ii)}$ $\delta E_{(iii)}$ $\Gamma^{s}_{(i)}$ $\Gamma^{s}_{(ii)}$ $\Gamma^{s}_{(iii)}$ $\delta E^{s}_{(i)}$ $\delta E^{s}_{(ii)}$ $\delta E^{s}_{(iii)}$
$1s_{0}$ $4554$ $7$ $<1$ $5$ $58$ $31$ $37$ $19$ $4$ $14$ $45$ $27$ $30$
$2s_{0}$ $4996$ $2$ $<1$ $1$ $33$ $15$ $20$ $5$ $<1$ $3$ $23$ $11$ $14$
$3s_{0}$ $5388$ $<1$ $1$ $<1$ $21$ $8$ $12$ $1$ $<1$ $<1$ $14$ $5$ $8$
$4s_{0}$ $5747$ $<1$ $1$ $<1$ $15$ $6$ $8$ $<1$ $<1$ $<1$ $9$ $3$ $5$
$5s_{0}$ $6081$ $<1$ $2$ $<1$ $11$ $4$ $6$ $<1$ $<1$ $<1$ $6$ $2$ $4$
$1(s\backslash d)_{1}$ $4110$ $36$ $28$ $24$ $-29$ $-13$ $-17$ n/a n/a n/a $-9$ $-5$ $-5$
$2(s\backslash d)_{1}$ $4462$ $49$ $23$ $29$ $-9$ $-1$ $-5$ $20$ $11$ $13$ $-10$ $-4$ $-6$
$3(s\backslash d)_{1}$ $4765$ $6$ $<1$ $3$ $20$ $14$ $13$ $8$ $2$ $5$ $14$ $11$ $9$
$4(s\backslash d)_{1}$ $4814$ $28$ $14$ $17$ $-8$ $-1$ $-4$ $11$ $6$ $7$ $-6$ $-2$ $-4$
$5(s\backslash d)_{1}$ $5130$ $23$ $7$ $13$ $5$ $3$ $3$ $14$ $5$ $8$ $2$ $1$ $1$
$1(p\backslash f)_{2}$ $4321$ $24$ $22$ $16$ $-8$ $-3$ $-5$ $5$ $7$ $4$ $-6$ $-5$ $-4$
$2(p\backslash f)_{2}$ $4658$ $30$ $19$ $19$ $-4$ $<1$ $-2$ $14$ $10$ $9$ $-5$ $-2$ $-3$
$3(p\backslash f)_{2}$ $4965$ $10$ $4$ $6$ $9$ $8$ $6$ $8$ $3$ $5$ $5$ $5$ $3$
$4(p\backslash f)_{2}$ $5001$ $16$ $11$ $10$ $-4$ $<1$ $-2$ $6$ $5$ $4$ $-3$ $<1$ $-2$
$5(p\backslash f)_{2}$ $5297$ $20$ $9$ $12$ $2$ $2$ $1$ $12$ $6$ $7$ $<1$ $<1$ $<1$
@@
Table 4: Contributions of $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}$ to the widths and masses of hybrid charmonium states with $\Sigma_{u}^{-}$ component, corresponding to $H_{3}[s_{0}]$, $H_{1}[(s\backslash d)_{1}]$ and $H_{4}[(p\backslash f)_{2}]$ multiplets of table 3. All dimension-full entries are in MeV units. The angular momentum of the heavy meson-antimeson pair is $l^{\prime}=\ell$. The subscript with Roman numbers refers to the model potential used (see Fig. 2). The superscript $s$ indicates the that the quantities refer to the heavy-meson pairs with strangeness. Energy contributions labeled as $<1$ should be understood as the absolute value of the contribution being smaller than $1$ MeV. The third column corresponds to $M^{(0)}=E^{(0)}+2m_{Q}$.
$n\,l_{\ell}$ $M_{b}^{(0)}$ $\Gamma_{(i)}$ $\Gamma_{(ii)}$ $\Gamma_{(iii)}$ $\delta E_{(i)}$ $\delta E_{(ii)}$ $\delta E_{(iii)}$ $\Gamma^{s}_{(i)}$ $\Gamma^{s}_{(ii)}$ $\Gamma^{s}_{(iii)}$ $\delta E^{s}_{(i)}$ $\delta E^{s}_{(ii)}$ $\delta E^{s}_{(iii)}$
$1s_{0}$ $11077$ $2$ $5$ $<1$ $107$ $51$ $71$ $3$ $<1$ $6$ $87$ $46$ $62$
$2s_{0}$ $11361$ $1$ $5$ $<1$ $67$ $28$ $42$ $4$ $<1$ $4$ $54$ $23$ $35$
$3s_{0}$ $11616$ $<1$ $5$ $<1$ $46$ $17$ $28$ $2$ $<1$ $2$ $36$ $14$ $22$
$4s_{0}$ $11851$ $1$ $4$ $<1$ $33$ $12$ $19$ $1$ $<1$ $<1$ $25$ $9$ $15$
$5s_{0}$ $12071$ $1$ $4$ $<1$ $25$ $9$ $15$ $<1$ $<1$ $<1$ $19$ $6$ $11$
$1(s\backslash d)_{1}$ $10783$ $81$ $55$ $64$ $-59$ $-14$ $-28$ n/a n/a n/a $-15$ $-7$ $-10$
$2(s\backslash d)_{1}$ $10980$ $129$ $45$ $74$ $-21$ $<1$ $-8$ $36$ $18$ $25$ $-29$ $-9$ $-15$
$3(s\backslash d)_{1}$ $11176$ $71$ $18$ $41$ $34$ $15$ $20$ $58$ $18$ $35$ $15$ $9$ $9$
$4(s\backslash d)_{1}$ $11219$ $14$ $11$ $8$ $25$ $21$ $17$ $<1$ $<1$ $<1$ $26$ $18$ $18$
$5(s\backslash d)_{1}$ $11380$ $63$ $17$ $35$ $6$ $3$ $4$ $36$ $11$ $20$ $-3$ $<1$ $-2$
$1(p\backslash f)_{2}$ $10895$ $63$ $39$ $40$ $-19$ $-4$ $-10$ $4$ $6$ $4$ $-14$ $-8$ $-9$
$2(p\backslash f)_{2}$ $11092$ $87$ $40$ $52$ $-13$ $<1$ $-6$ $29$ $17$ $19$ $-17$ $-7$ $-10$
$3(p\backslash f)_{2}$ $11284$ $75$ $27$ $43$ $9$ $7$ $6$ $44$ $18$ $26$ $-2$ $1$ $<1$
$4(p\backslash f)_{2}$ $11346$ $3$ $4$ $2$ $20$ $18$ $13$ $<1$ $<1$ $<1$ $19$ $14$ $12$
$5(p\backslash f)_{2}$ $11480$ $64$ $23$ $37$ $<1$ $2$ $<1$ $32$ $13$ $19$ $-6$ $-1$ $-3$
@@
Table 5: Contributions of $B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)}$ to the widths and masses of hybrid bottomonium states with $\Sigma_{u}^{-}$ component, corresponding to $H_{3}[s_{0}]$, $H_{1}[(s\backslash d)_{1}]$ and $H_{4}[(p\backslash f)_{2}]$ multiplets of table 3. All dimension-full entries are in MeV units. The angular momentum of the heavy meson-antimeson pair is $l^{\prime}=\ell$. The subscript with Roman numbers refers to the model potential used (see Fig. 2). The superscript $s$ indicates that the quantities refer to the heavy-meson pairs with strangeness. Energy contributions labeled as $<1$ should be understood as the absolute value of the contribution being smaller than $1$ MeV. The third column corresponds to $M^{(0)}=E^{(0)}+2m_{Q}$.
@@
We display our results for the energy shifts caused by the heavy-meson pairs to the hybrid masses and the decay widths of the hybrid states into the heavy-meson pairs in tables 4 and 5 for charmonium and bottomonium hybrids, respectively. In the charmonium sector we find that the energy contributions are, with a few exceptions, below the $30$ MeV, thus being smaller than the $\mathcal{O}(1/m_{Q})$ spin-dependent operator contributions Brambilla:2018pyn ; Brambilla:2019jfi ; Soto:2023lbh and also smaller than the expected size of the spin-independent ones, which so far have not been computed. The energy contributions tend to decrease with increasing $n$ and $\ell$. The widths range between $1$ to $20$ MeV per channel [^3] Notice, that neutral and charged $D$ meson contributions are added up in columns 4 to 9 of table 4.. The larger widths are found in states closer to the heavy-meson pair thresholds, with the values being very sensitive to the energy gap between the state and threshold. Adding up the widths for the two channels studied, we find values that range from slightly larger than the semi-inclusive widths for transitions to quarkonium Oncala:2017hop ; Brambilla:2022hhi to much smaller ones. The latter, however, correspond to higher lying hybrid charmonium states that can also decay to higher mass heavy-meson pairs not considered here.
@@
In the bottomonium sector, the energy contributions are larger than in the charmonium sector, with up to ten states having contributions larger than $30$ MeV. As a consequence for bottomonium hybrids the $B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)}$ contribution to the mass is significantly larger than the spin-dependent one. The widths are a factor of $2$ to $3$ larger than in the charmonium sector. The larger contributions in the bottomonium sector can be traced back to the fact that the coupling operator is not heavy-quark mass suppressed and the less extended hybrid bottomonium wave functions having a larger overlap with the coupling potential than the more extended wave functions of hybrid charmonia. In contrast, semi-inclusive transition widths to quarkonium are smaller for bottomonium hybrids than for charmonium hybrids Oncala:2017hop ; TarrusCastella:2021pld ; Brambilla:2022hhi and consequently also smaller than decay widths to $B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)}$. As in the charmonium sector, the lowest mass $s_{0}$ states have the largest contributions to the mass, while the larger widths are found in the lowest lying $(s\backslash d)_{1}$ states.
@@
It is perhaps unexpected that the widths for $s_{0}$ are smaller than those of other states of similar mass, while the energy shifts are larger. This can be explained as follows. The transition form factors into $S$-wave heavy-meson pairs, $a_{n0;S}(k)$, have the largest peak, resulting in large values for the energy shifts produced by the heavy-meson pairs, however, they decrease very rapidly to zero as $k$ increases. As the $s_{0}$ states are far away from the thresholds the values of $p_{r}$ are large enough for the widths to correspond to the region where the form factors are small. Moreover, as the energy contribution is overall positive, it tends to increase the value of hybrid mass and, correspondingly, of $p_{r}$, further decreasing the value of the width.
@@
In table 6 we show the spectra of the pure $\Pi_{u}$ states that do not couple with the two lowest lying thresholds at leading order. The lowest mass states ($1p_{1}$), in both sectors, are interesting since the decays into $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ are $1/m_{Q}$-suppressed and are below any other heavy-meson pair threshold. Furthermore, two of the spin-triplet states of corresponding multiplet have explicitly exotic $j^{pc}$. If they could be measured, identifying these states as hybrids should be straightforward due to these two very distinct characteristics.
$n\,l_{\ell}$ $M^{(0)}_{c}$ $M^{(0)}_{b}$
$1p_{1}$ $4238$ $10840$
$2p_{1}$ $4610$ $11059$
$3p_{1}$ $4954$ $11267$
$4p_{1}$ $5277$ $11466$
$5p_{1}$ $5583$ $11657$
$1d_{2}$ $4429$ $10945$
$2d_{2}$ $4786$ $11160$
$3d_{2}$ $5118$ $11364$
$4d_{2}$ $5433$ $11560$
$5d_{2}$ $5732$ $11748$
@@
Table 6: Masses of the pure $\Pi_{u}$ hybrid quarkonium states that do not couple to $D_{(s)}^{(*)}\bar{D}_{(s)}^{(*)}(B_{(s)}^{(*)}\bar{B}_{(s)}^{(*)})$ corresponding to the $H_{2}[p_{1}]$ and $H_{5}[d_{2}]$ multiplets of table 3. The masses are obtained as $M^{(0)}=E^{(0)}+2m_{Q}$ and are given in MeV units.
@@
Figure 3: Comparison of the mass and widths of hybrid charmonium states in tables 4 and 6 with experimental states above threshold listed in the PDG Workman:2022ynf . The masses and widths of the hybrid states are obtained as the average of the results corresponding to the three model potentials in Fig. 2. The bars correspond to the sum in quadrature of the standard deviation of the results for the three models and the size of higher order corrections $\mathcal{O}(\Lambda_{\rm QCD}/m_{c})$.